OpenStudy (anonymous):

How long is line segment . Select one of the options below as your answer: A. 4 feet B. 6 feet C. 8 feet D. 10 feet

OpenStudy (anonymous):

ad equals 9

OpenStudy (unklerhaukus):

and you are looking to find \(\overline{DB}\) right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[\overline{AD}:\overline{CE}=\overline{DB}:\overline{BE}\]

OpenStudy (anonymous):

the number on one of the anweres

OpenStudy (unklerhaukus):

\[\frac{\overline{AD}}{\overline{CE}}=\frac{\overline{DB}}{\overline{BE}}\] \[{\overline{BE}\frac{\overline{AD}}{\overline{CE}}=\overline{DB}}\] \[\overline{DB}=8[\text {ft}]\frac{9[\text {ft}]}{12[\text {ft}]}\]

OpenStudy (anonymous):

aight

OpenStudy (unklerhaukus):

am i making sense?